# Communication-Efficient and Tensorized Federated Fine-Tuning of Large Language Models

### v1

@@ 4 Proposed Algorithm
In this section, we introduce the proposed FedTT method, which is built on a novel parameter-efficient tensorized adapter. This adapter is integrated into clients’ local models to effectively adapt LLMs. We begin with discussing the structure and setup of the tensorized adapters, followed by the introduction of FedTT and its enhanced version, FedTT+, which can be selected based on communication budget constraints. The overall workflow of our FedTT method is shown in Fig. 1 (b).
@@ 4.1 Tensorized Adapters
To facilitate fine-tuning in our proposed tensorized FL framework, we incorporate tensorized adapters designed to adapt LLMs with minimal additional parameters. These adapters are underpinned by the novel tensorized linear layer [39, 42], which replaces the matrix weight in linear layers with more parameter-efficient TT format weights. This section first details the architecture of the tensorized linear layer and then elucidates the process of constructing tensorized adapters using this innovative layer.
In Sec. 3.2, we introduced how to represent a weight matrix in the TT format, where the TT format tensor factors $\mathcal{G}_{1},\ldots,\mathcal{G}_{J}$ are initialized, stored, and updated during the fine-tuning process. The tensorized layer is successfully integrated into matrix-based LLMs, performing tensor contraction over the list of tensor factors and reshaping them to match the traditional weight matrix $\boldsymbol{W}$. Compared to traditional matrix-based linear layers, the tensorized layers only store the smaller tensor factors, significantly reducing the number of parameters. For instance, a weight matrix of size $16\times 768$, typically containing around $6.1K$ trainable parameters, can be represented by tensor factors totaling approximately $0.6K$ trainable parameters. This dramatic reduction is achieved by assuming a tensor rank of 5 and reshaping $\mathcal{W}$ into dimensions $[4,4,8,8,12]$.
We now introduce the tensorized adapters, built based on the tensorized layers shown in Fig. 1. These adapters utilize a bottleneck structure similar to that described in the original sequential adapter methodology [7], comprising two tensorized linear layers with a nonlinear layer in between. The bottleneck structure further reduces the number of trainable parameters by decreasing the channels connected between the two tensorized layers and the nonlinearity layer, from the size of the hidden dimension to a smaller number, such as 64.
As illustrated in Fig. 1 (b), the tensorized adapter is strategically placed after the attention and MLP components of an encoder/decoder block. In practical applications, we compress the classification head (linear layer) into a tensorized layer and make it trainable for sequence classification tasks, as shown in Fig. 1 (c). Instead, we retain the original language model head without compression for general language modeling tasks, as experiments show a significant reduction in performance when the language modeling head is compressed.
@@ 4.2 FedTT Method
During communication round $t$, each client $c_{i}\in\mathcal{C}$ updates its trainable parameters $\bm{w}^{(t)}_{i}$, including the tensorized classifier, tensorized adapter, and bias terms, for $K$ local updates. Afterward, each client sends their updated parameters, $\bm{w}^{(t)+K}_{i}$, to the central server. The server then aggregates the weights from all clients as $\bm{w}^{(t+1)}=\sum_{i=1}^{N}\bm{w}^{(t)+K}_{i}$ to obtain the updated global weights for the next training round and sends the aggregated weights $\bm{w}^{(t+1)}$ back to the clients. Note that while we have primarily described cross-silo FL scenarios, the large-scale cross-device FL setting operates similarly, with the key distinction being that in each communication round, the server sends the aggregated weights to a randomly chosen subset of clients $\mathcal{S}\subseteq\mathcal{C}$, rather than to all clients in $\mathcal{C}$.
Unlike most previous Federated PEFT methods such as LoRA, which require transmitting large matrices between clients and the central server, FedTT significantly reduces communication overhead by only transferring small tensor factors. This approach leads to over 10×10$\times$10 × less communication cost compared to LoRA adapters, as demonstrated in Sec. 4.1, making FedTT particularly advantageous in FL scenarios where efficient communication is critical.
@@ 4.3 FedTT+ Method
$\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}_{i,1}\right)\times_{3,1}\ldots\times_{3,1}\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}_{i,J}\right)\neq\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}_{i,1}\times_{3,1}\ldots\times_{3,1}\mathcal{G}_{i,J},$ (2)
where $\mathcal{G}_{i,j}$ is the tensor factor $j$ for clientuser $c_{i}$.
$\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}^{(t)}_{i,1}\right)\times_{3,1}\mathcal{G}^{(t-1)}_{2}\times_{3,1}\ldots\times_{3,1}\mathcal{G}_{J}^{(t-1)}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathcal{G}^{(t)}_{i,1}\times_{3,1}\mathcal{G}^{(t-1)}_{2}\times_{3,1}\ldots\times_{3,1}\mathcal{G}_{J}^{(t-1)}\right).$
This modification to FedTT can enhance accuracy in FL scenarios, particularly in cases of severe data heterogeneity. The detailed algorithm is presented in Algorithm 2.
FedTT+ operates similarly to FedTT, but with a key difference: in each communication round $t$, an index $r$ is selected from the set $r\in\{2,\ldots,J-1\}$ (line 3 in Alg. 2). In this process, $\{\mathcal{G}_{1},\mathcal{G}_{r},\mathcal{G}_{J}\}$ are set as trainable parameters (as the first and last tensors are always trained), while the other factors, $\{\mathcal{G}_{2},\ldots,\mathcal{G}_{r-1},\mathcal{G}_{r+1},\ldots,\mathcal{G}_{J-1}\}$, remain frozen (line 5 and 6 in Alg. 2). As a result, clients only send their updated $\{\mathcal{G}_{i,1},\mathcal{G}_{i,r},\mathcal{G}_{i,J}\}$ to the server for aggregation, significantly reducing communication overhead.
[36, 37] have also demonstrated that freezing certain parameters improves the performance of LoRA in the presence of data heterogeneity, particularly with larger models like RoBERTa-large. In the numerical section, we compare FedTT+ with their method under the same settings and show that our approach achieves comparable or better accuracy while using even fewer trainable parameters.
@@
3:         $r\leftarrow$ (choose index $r$ s.t. mod$(t,{\scriptsize J-2})$ $=$ $r-1$)
@@
5 Empirical Results
5.1 Experiments Setup
We conducted extensive experiments to evaluate the performance of the proposed algorithms across various language models. For the BERT-family models, we utilized RoBERTa-base [43], DeBERTa-base [44], and RoBERTa-large [43], while for large-scale models, we employed LLaMA-2 [2] (result for LLaMA-2 is presented in Appendix B). Using these models, we compared the proposed algorithm against several PEFT methods in FL scenarios, including BitFit, LoRA, Adapter, Prefix-Tuning, and Prompt-Tuning. Additionally, we benchmarked it against SOTA federated PEFT methods such as FFA-LoRA [36] and RoLoRA [37].
In our experiments, we consider two main FL scenarios: the cross-silo FL scenario [45] and the large-scale cross-device FL scenario [46]. The cross-silo scenario is suitable for networks with typically fewer than $100$ clients. In this case, the server sends the updated model to all clients, meaning every client participates in training during each communication round. In contrast, large-scale cross-device FL is more appropriate for environments with thousands of clients, where only a randomly selected subset of clients is involved in each training round. For cross-silo FL, the number of clients is set to $n\in\{5,10,20,50\}$, while for large-scale cross-device FL, similar to [14], we randomly select $10$ clients from a pool of $1000$. All experiments were conducted using the AdamW optimizer [47], with a similar learning rate and batch size setup across different methods. We performed the experiments on NVIDIA A6000 and V100 GPUs.
@@ 5.1 Experiments Setup
Table 2: Comparative analysis of various federated PEFT methods using the RoBERTa-Base model, focusing on cross-silo FL for MRPC and SST-2 tasks and large-scale cross-device FL for other tasks. The reported accuracy for federated LoRA, Adapter, Prefix, and BitFit methods is sourced from [14].
@@ 5.2 Performance on the BERT Family
We conducted experiments using the Generalized Language Understanding Evaluation (GLUE) benchmark [48], employing the complete training dataset for each task. We recorded the best validation results after the $100$ communication round for cross-silo FL and after the $1000$ communication round for large-scale cross-device FL. Initially, we compared FedTT and FedTT+ with various PEFT methods in a cross-silo FL setting, using the DeBERTa-base model and assuming independent and identically distributed (i.i.d.) data among clients. We set the number of clients to $5$ and the local training epochs to $1$. The results, shown in Table 1, indicate that FedTT achieves a higher average score compared to other methods with lower than $0.15M$ trainable parameters. Notably, FedTT exhibits only a $0.5\%$ performance gap compared to LoRA with a rank of $8$, which has 5×5$\times$5 × more trainable parameters. FedTT+ achieves comparable accuracy to other methods while drastically reducing the number of trainable parameters. It exhibits only a $0.6\%$ performance gap compared to LoRA with a rank of $4$, while using 6×6$\times$6 × fewer trainable parameters, which significantly lowers communication overhead.
We then used the RoBERTa-base model to further evaluate FedTT against other PEFT methods. For the MRPC and SST-2 tasks, we employed a cross-silo FL setup with $10$ clients, while for the other tasks, we applied a large-scale cross-device FL configuration with $1000$ clients, randomly selecting $10$ clients per round. Our settings for the RoBERTa-base model align with those in [14], allowing us to leverage their results of PEFT methods like LoRA, Adapter, Prefix, and BitFit under FL settings. The results in Table 2 show that FedTT achieves a higher average score with the fewest trainable parameters among all methods. Additional details regarding Tables 1 and 2 are provided in Appendices A.3 and A.4, respectively.
5.3 Impact of Data Heterogeneity
Similar to [37], we simulate varying levels of data heterogeneity by using different numbers of clients: $3$ clients (i.i.d. data distribution), $20$ clients (mild heterogeneity), and $50$ clients (severe heterogeneity), with $20$ local updates to further increase data heterogeneity. Additional details are provided in Appendix A.5. Our RoBERTa-Large model settings align with those in [49], allowing us to leverage their results for LoRA, FFA-LoRA, and RoLoRA methods. As shown in Table 3, FedTT+ consistently outperforms these methods across different heterogeneity settings while utilizing fewer parameters. Fig. 2 illustrates test accuracy versus the number of communication rounds in the severe heterogeneity scenario, highlighting the faster convergence of FedTT+ compared to RoLoRA across various tasks.
Table 3: Comparison of SOTA cross-silo FL methods using RoBERTa-Large models under varying degrees of data heterogeneity. The accuracies for LoRA, FFA-LoRA, and RoLoRA methods are sourced from [37].
@@ 5.3 Impact of Data Heterogeneity
RoBERTa-Large (FedTT+) $\bm{27K}$ $\bm{95.64}$ $\bm{94.05}$ $\bm{87.99}$ $\bm{87.27}$ $\bm{91.24}$
@@
RoBERTa-Large (FedTT+) $\bm{27K}$ $\bm{95.64}$ $\bm{92.60}$ $\bm{87.46}$ $\bm{87.69}$ $\bm{90.85}$
@@
RoBERTa-Large (FedTT+) $\bm{27K}$ $94.50$ $\bm{90.17}$ $\bm{86.65}$ $\bm{86.28}$ $\bm{89.40}$

### v2

@@ 4 Proposed Algorithm
In this section, we introduce the proposed FedTT method, which is built on a novel parameter-efficient tensorized adapter. This adapter is integrated into clients’ local models to effectively adapt LLMs. We begin with discussing the structure and setup of the tensorized adapters, followed by the introduction of FedTT and its enhanced version, FedTT+, which can be selected based on communication budget constraints. The workflow of our FedTT method is shown in Fig. 1 (b).
@@ 4.1 Tensorized Adapters
To facilitate fine-tuning in a FL framework, we incorporate tensorized adapters designed to adapt LLMs with minimal additional parameters. These adapters are underpinned by the novel tensorized linear layer Yang et al. (2024b, a), which replaces the matrix weight in linear layers with more parameter-efficient TT format weights. This section first details the architecture of the tensorized linear layer and then elucidates the process of constructing tensorized adapters.
In Sec. 3.2, we introduced how to represent a weight matrix in the TT format, where the TT format tensor factors $\mathcal{G}_{1},\ldots,\mathcal{G}_{J}$ are initialized, stored, and updated during the fine-tuning process. The tensorized layer is successfully integrated into matrix-based LLMs, performing tensor contraction over the list of tensor factors and reshaping them to match the traditional weight matrix $\bm{W}$. Compared to traditional matrix-based linear layers, the tensorized layers only store the smaller tensor factors, significantly reducing the number of parameters.
We now introduce the tensorized adapters, built based on the tensorized layers in Fig. 1. These adapters use a bottleneck structure similar to that in the original sequential adapter methodology Houlsby et al. (2019), comprising two tensorized linear layers with a nonlinear layer in between. The bottleneck structure further reduces the number of trainable parameters by decreasing the channels connected between the two tensorized layers and the nonlinearity layer, from the size of the hidden dimension to a smaller number, such as 64. As illustrated in Fig. 1 (b), the tensorized adapter is strategically placed after the attention and MLP components of an encoder/decoder block. In practical applications, we compress the classification head (linear layer) into a tensorized layer and make it trainable for sequence classification tasks, as shown in Fig. 1 (c). Instead, we retain the original language model head without compression for general language modeling tasks, as experiments show a significant reduction in performance when the language modeling head is compressed.
@@ 4.2 FedTT Method
During the communication round $t$, each client $c_{i}\in\mathcal{C}$ updates its trainable parameters $\bm{w}^{(t)}_{i}$, including the tensorized classifier, tensorized adapter, and bias terms (if they exist), for $K$ local updates. Afterward, each client sends its updated parameters, $\bm{w}^{(t)+K}_{i}$, to the central server. The server then aggregates the weights as $\bm{w}^{(t+1)}=1/N\sum_{i=1}^{N}\bm{w}^{(t)+K}_{i}$ for the next training round and sends updated weights $\bm{w}^{(t+1)}$ back to the clients. Note that while we have primarily described cross-silo FL scenarios, the LSCD FL setting operates similarly, with the key distinction being that in each communication round, the server sends the aggregated weights to a randomly chosen subset of clients $\mathcal{S}\subseteq\mathcal{C}$, rather than to all clients in $\mathcal{C}$.
Compared with most previous federated PEFT methods, FedTT significantly reduces communication overhead by only transferring small tensor factors. This approach leads to over 10×10$\times$10 × communication reduction compared to LoRA adapters, as demonstrated in Sec. 4.1, making FedTT particularly advantageous in FL scenarios where efficient communication is critical.
@@ 4.3 FedTT+ Method
3:         $r\leftarrow$ (choose index $r$ s.t. mod$(t,{J-2})$ $=$ $r-1$)
@@
$\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}_{i,1}\right)$$\times_{3,1}\ldots\times_{3,1}\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}_{i,J}\right)$ (2)
$\neq\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}_{i,1}\times_{3,1}\ldots\times_{3,1}\mathcal{G}_{i,J},$
where $\mathcal{G}_{i,j}$ is the tensor factor $j$ for clientuser $c_{i}$.
@@
$\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{G}^{(t)}_{i,1}\right)\times_{3,1}\mathcal{G}^{(t-1)}_{2}\times_{3,1}\ldots\times_{3,1}\mathcal{G}_{J}^{(t-1)}$
$=\frac{1}{N}\sum_{i=1}^{N}\left(\mathcal{G}^{(t)}_{i,1}\times_{3,1}\mathcal{G}^{(t-1)}_{2}\times_{3,1}\ldots\times_{3,1}\mathcal{G}_{J}^{(t-1)}\right).$
This modification to FedTT can improve accuracy, particularly in cases of severe data heterogeneity. The detailed algorithm is presented in Algorithm 2.
FedTT+ operates similarly to FedTT, but with a key difference: in each communication round $t$, an index $r$ is selected from the set $r\in\{2,\ldots,J-1\}$ (line 3 in Alg. 2). In this process, $\{\mathcal{G}_{1},\mathcal{G}_{r},\mathcal{G}_{J}\}$ are set as trainable parameters (as the first and last tensors are always trained), while the other factors, $\{\mathcal{G}_{2},\ldots,\mathcal{G}_{r-1},\mathcal{G}_{r+1},\ldots,\mathcal{G}_{J-1}\}$, remain frozen (line 5 and 6 in Alg. 2). As a result, clients only send their updated $\{\mathcal{G}_{i,1},\mathcal{G}_{i,r},\mathcal{G}_{i,J}\}$ to the server for aggregation, significantly reducing communication overhead. Note that the classification head is always trainable and present in all communication rounds.
Sun et al. (2024); Chen et al. have also demonstrated that freezing certain parameters improves the performance of LoRA in the presence of data heterogeneity, particularly with larger models like RoBERTa-large. In the numerical section, we compare FedTT+ with their method under the same settings and show that our approach achieves comparable or better accuracy while using even fewer trainable parameters.
5 Numerical Results
5.1 Experiments Setup
Table 2: Comparative analysis of various federated PEFT methods using the RoBERTa-Base model. The reported accuracy for federated LoRA, Adapter, Prefix, and BitFit methods is sourced from Zhang et al. (2023).
@@ 5.1 Experiments Setup
We conduct extensive experiments to evaluate the performance of the proposed algorithms across various language models. For the BERT-family models, we utiliz RoBERTa-base Liu et al. (2019), DeBERTa-base He et al. (2020), and RoBERTa-large Liu et al. (2019), while for large-scale models, we employed LLaMA-2 Touvron et al. (2023). Using these models, we compare the proposed algorithm against several PEFT methods in FL scenarios, including BitFit, LoRA, Adapter, Prefix-Tuning, and Prompt-Tuning. Additionally, we benchmark it against SOTA federated PEFT methods such as FFA-LoRA Sun et al. (2024) and RoLoRA Chen et al. .
We consider two main FL scenarios: the cross-silo FL scenario Kairouz et al. (2021) and the LSCD FL scenario Lai et al. (2022). The cross-silo scenario is suitable for networks with typically fewer than $100$ clients. In this case, the server sends the updated model to all clients, meaning every client participates in training during each communication round. In contrast, LSCD FL is more appropriate for environments with thousands of clients, where only a randomly selected subset of clients is involved in each training round. For cross-silo FL, the number of clients is set to $n\in\{5,10,20,50\}$, while for LSCD FL, similar to Zhang et al. (2023), we randomly select $10$ clients from a pool of $1000$. All experiments are conducted using the AdamW optimizer Loshchilov and Hutter (2018), with a similar learning rate and batch size setup across different methods. We perform the experiments on NVIDIA A6000 and V100 GPUs.
@@ 5.2 Performance on the BERT Family
Table 3: Comparison of SOTA cross-silo FL methods using RoBERTa-Large models under varying degrees of data heterogeneity. The accuracies for LoRA, FFA-LoRA, and RoLoRA methods are sourced from Chen et al. .
@@
RoBERTa-Large (FedTT) $51K$ $94.38$ $93.01$ $88.30$ $87.20$ $90.72$
RoBERTa-Large (FedTT+) $\bm{28K}$ $\bm{95.64}$ $\bm{94.05}$ $\bm{88.99}$ $\bm{88.27}$ $\bm{91.74}$
@@
RoBERTa-Large (FedTT) $51K$ $94.15$ $91.38$ $86.25$ $86.53$ $89.58$
RoBERTa-Large (FedTT+) $\bm{28K}$ $\bm{95.64}$ $\bm{92.60}$ $\bm{87.76}$ $\bm{88.11}$ $\bm{91.03}$
@@
RoBERTa-Large (FedTT) $51K$ $94.38$ $\bm{90.55}$ $85.47$ $85.27$ $88.92$
RoBERTa-Large (FedTT+) $\bm{28K}$ $94.50$ $90.17$ $\bm{86.65}$ $\bm{86.28}$ $\bm{89.40}$
Table 4: Comparison of cross-silo FL methods using RoBERTa-Large models under varying differential privacy guarantees. The accuracies for LoRA and FFA-LoRA methods are sourced from Sun et al. (2024).
Priv. Budget Method # Param. SST-2 QNLI QQP MNLI
$\epsilon=6$ LoRAr=8 $1.57M$ $93.70$ $84.99$ $82.11$ $39.46$
FFA-LoRAr=8 $0.79M$ $93.73$ $87.27$ $83.31$ $78.81$
FedTT $\bm{0.24M}$ $\bm{93.80}$ $\bm{87.43}$ $\bm{84.61}$ $\bm{85.45}$
$\epsilon=3$ LoRAr=8 $1.57M$ $93.32$ $83.94$ $82.08$ $35.82$
FFA-LoRAr=8 $0.79M$ $93.59$ $86.18$ $83.03$ $77.42$
FedTT $\bm{0.24M}$ $\bm{93.96}$ $\bm{86.64}$ $\bm{84.36}$ $\bm{85.08}$
$\epsilon=1$ LoRAr=8 $1.57M$ $94.32$ $88.95$ $81.28$ $33.80$
FFA-LoRAr=8 $0.79M$ $94.32$ $90.35$ $82.50$ $75.05$
FedTT $\bm{0.24M}$ $\bm{95.64}$ $\bm{91.67}$ $\bm{83.11}$ $\bm{83.81}$
We conduct experiments using the Generalized Language Understanding Evaluation (GLUE) benchmark Wang et al. (2018), employing the complete training dataset for each task. We record the best validation results after the $100$ communication round for cross-silo FL and after the $1000$ communication round for LSCD FL. Initially, we compare FedTT and FedTT+ with various PEFT methods in a cross-silo FL setting, using the DeBERTa-base model and assuming independent and identically distributed (i.i.d.) data among clients. We set the number of clients to $5$ and the local training epochs to $1$. The results, shown in Table 1, indicate that FedTT achieves a higher average score compared to other methods with lower than $0.15M$ trainable parameters. Notably, FedTT exhibits only a $0.5\%$ performance gap compared to LoRA with a rank of $8$, which has 5×5$\times$5 × more trainable parameters. FedTT+ achieves comparable accuracy to other methods while drastically reducing the number of trainable parameters. It exhibits only a $0.6\%$ performance gap compared to LoRA with a rank of $4$, while using 6×6$\times$6 × fewer trainable parameters, which significantly lowers communication overhead.
We then use the RoBERTa-base model to further evaluate FedTT against other PEFT methods. For the MRPC and SST-2 tasks, we employ a cross-silo FL setup with $10$ clients, while for the other tasks, we apply a LSCD FL configuration with $1000$ clients, randomly selecting $10$ clients per round. Our settings for the RoBERTa-base model align with those in Zhang et al. (2023), allowing us to leverage their results of PEFT methods like LoRA, Adapter, Prefix, and BitFit under FL settings. Table 2 shows that FedTT achieves a higher average score with the fewest trainable parameters among all methods. Additional details regarding Tables 1 and 2 are provided in Appendices A.3 and A.4, respectively.
5.3 Impact of Data Heterogeneity
@@
Fed. Set. Model & Method # Param. COPA ReCoRD SQuAD Avg.
large-scale FL LLaMA2-7B (LoRAr=8) $4.19M$ $87$ $81.0$ $90.49$ $86.16$
LLaMA2-7B (FedTT) $0.52M$ $90$ $80.1$ $89.32$ $86.47$
cross-silo FL LLaMA2-13B (LoRAr=8) $6.55M$ $89$ $83.6$ $91.07$ $87.89$
LLaMA2-13B (FedTT) $0.64M$ $89$ $83.5$ $90.06$ $87.52$
LLaMA2-13B (FedTT+) $0.18M$ $88$ $83.7$ $90.40$ $87.70$
Table 6: Total transmitted messages for the MNLI task.
Methods Total transmitted messages (KB) Comm. overhead LoRA 111172    height 0.28cm depth 3pt width 1.5pt $\times$1.881 BitFit 11390      height 0.28cm depth 3pt width 1.5pt $\times$0.621 RoLoRA 11624      height 0.28cm depth 3pt width 1.5pt $\times$1.001 Prompt 111523      height 0.28cm depth 3pt width 1.5pt $\times$2.441 FedTT 11234      height 0.28cm depth 3pt width 1.5pt $\times$0.371 FedTT+ 1178        height 0.28cm depth 3pt width 1.5pt $\times$0.121
Similar to Chen et al. , we simulate varying levels of data heterogeneity by using different numbers of clients: $3$ clients (i.i.d. data distribution), $20$ clients (mild heterogeneity), and $50$ clients (severe heterogeneity), with $20$ local updates to further increase data heterogeneity. To ensure a comparable number of trainable parameters, we make the last $6$ layers trainable for the SST-2 dataset and the last $9$ layers for the other datasets in LoRA, FFA-LoRA, RoLoRA, and FedTT. In contrast, for FedTT+, all layers are made trainable. Additional details are provided in Appendix A.5. Our RoBERTa-Large model settings align with those in Che et al. (2023), allowing us to leverage their results for LoRA, FFA-LoRA, and RoLoRA methods. As shown in Table 3, FedTT+ consistently outperforms the other methods across varying heterogeneity settings while utilizing fewer parameters. This further demonstrates that, under a comparable number of trainable parameters and in the presence of data heterogeneity, FedTT+ outperforms FedTT, highlighting the effectiveness of adaptively freezing tensorized adapters in addressing data heterogeneity. Fig. 2 illustrates test accuracy versus the number of communication rounds in the severe heterogeneity scenario, showing that FedTT+ achieves faster convergence than RoLoRA across multiple tasks. An additional experiment examining the effect of data heterogeneity using an alternative heterogeneity measure is provided in Appendix B.
5.4 Performance on Larger Models
We use SuperGLUE tasks Wang et al. (2019) and a generation task (SQuAD Rajpurkar et al. (2016)) to compare FedTT and FedTT+ with LoRA. The results are shown in Table 5. We use the LLaMA2-7B model to simulate LSCD FL with $1000$ clients, where $10$ of them are chosen randomly in each round for training. Additional details are provided in Appendix A.6. FedTT has performance nearly identical to LoRA with a rank of $8$, while achieving about 10×10$\times$10 × lower communication overhead. For cross-silo FL, we utilize the LLaMA2-13B model with $10$ clients. As seen in Table 5, both FedTT and FedTT+ demonstrate performance close to LoRA, with 10×10$\times$10 × and 30×30$\times$30 × lower communication overhead, respectively.
5.5 Communication Cost Analysis
In this section, we present the communication cost analysis for the MNLI task of Table 1 in Table 6. The analysis for all tasks of Table 1 and Table 3 can be found in Appendix C. Following Guo et al. (2024), we compare the proposed method with baselines in terms of: (i) Up-link message size (in KB) per communication round, (ii) Number of communication rounds required to reach $95$% of the prediction accuracy reported in Table 1, and (iii) Total size of transmitted messages. As demonstrated, our proposed algorithms effectively minimize both the up-link message size and the total size of transmitted messages.
5.6 Differential privacy guarantees
We compare our proposed FedTT algorithm with baseline methods under differential privacy guarantees. Definitions and theoretical results related to DP are provided in Appendix D. Following Sun et al. (2024), we train the RoBERTa-Large model with a learning rate of $31e-$ 3 and three clients. We implement the DP-SGD algorithm using the Opacus Yousefpour et al. (2021) library with privacy parameters $\delta=1e-5$ and three different privacy budgets, $\epsilon\in\{1,3,6\}$. The clipping threshold is chosen from $C\in\{2,5\}$.
We follow the experimental setup from Sun et al. (2024), enabling direct evaluation against LoRA and FFA-LoRA. The results in Table 4 demonstrates that FedTT consistently achieves higher accuracy across different privacy budgets while using fewer trainable parameters. Notably, the largest performance gap is observed in the MNLI task, a three-class classification problem.
5.7 Tensor Rank Analysis
In this section, we evaluate the performance of the proposed FedTT method with varying tensor ranks across four datasets using the DeBERTa-base model. The number of clients is set to $5$, each performing one local update per round. We use a learning rate of $31e-$ 3 and a batch size of $32$. The best validation results are reported after $20$ communication rounds. As shown in Table 7, higher tensor ranks lead to better average accuracy, albeit at the cost of an increased number of trainable parameters. In our experiments, we set the rank to $5$, as it offers a favorable trade-off between accuracy and model size.
Table 7: Tensor rank analysis using the DeBERTa-Base model.
Model & Method # Param. SST-2 QNLI QQP MNLI Avg.
FedTTr=2 $0.03M$ $93.58$ $90.76$ $86.79$ $85.29$ $89.10$
FedTTr=5 $0.06M$ $93.00$ $91.87$ $87.65$ $85.91$ $89.61$
FedTTr=10 $0.17M$ $93.69$ $92.48$ $88.22$ $87.01$ $90.35$